Live GMAT classes with Target Test Prep start Nov 21 and Nov 22. First class free.

See classes
Target Test Prep · GMAT

Choose how you want to prepare

Learn live with an expert or move at your own pace. Every option includes the complete TTP study system.

Newsweek Readers’ Choice Awards 2026 winner: Target Test Prep, #1 Test Prep

Target Test Prep was voted #1 Best Test Prep in the 2026 Newsweek Readers’ Choice™ Awards.

★★★★★5.0563 reviews
GMATLiveTeachStarts Nov 22
Josh Braslow, Target Test Prep GMAT expert
LIVE ONLINE CLASSES

Get Ready for GMAT Test Day Faster with Live Online Classes

with Josh Braslow, 20+ years in test prep

Nov 22 · Josh Braslow
Sun · 9:30 AM to 12:30 PM ET
Included
40 hours of live online classes + 6 months of TTP OnDemand
  • Attend the first class for free
  • Every class is recorded, so you never fall behind
View classes & enroll
Limited seats availableTarget Test Prep
EALiveTeachOnDemandStart anytime
EXECUTIVE ASSESSMENT

Target Test Prep EA OnDemand

Self-paced EA prep. Study on your schedule.

EXECUTIVE ASSESSMENT

with

165+ EA score guarantee
$05-day trial no automatic billing
Schedule
Included
40 hours of live online classes plus six months of access to the complete TTP EA OnDemand course.
  • 165+ EA Score Guarantee
  • 4,100+ Quant, Verbal, and Integrated Reasoning practice questions
  • 400+ hours of in-depth video lessons
  • 3,000+ step-by-step video solutions
View EA class & enroll Start free 5-day trial
Limited cohort · enrollment openTrial includes full course accessTarget Test Prep
GMATOnDemandStart anytime
SELF-PACED MASTERCLASS

Target Test Prep GMAT OnDemand

Complete access from day one. Study on your schedule.

715+ score guarantee
$0to start then $127/mo
  • Personalized study plan and analytics
  • Thousands of lessons and practice questions

Help with Probablility Question

Expert replies
by tim0thy » Fri Oct 16, 2009 11:14 am
It's a simple one but I'm not sure what the best approach is.

If there are six numbers numbered 1-6, what is the probability of choosing numbers 1, 2, and 3?


My approach: because order doesn't matter I used the combination 6 choose 3 which equals 20 possible combinations. Then there is only one combination out of these 20 that has 1, 2, and 3 so the probability is 1/20. Is that right?

If anyone can post with an explanation that would be fantastic.
Join the discussion
Source: — Problem Solving |

by uymba » Fri Oct 16, 2009 4:48 pm
the general rule for probablity is:

number of desired outcomes
---------------------------------
number of total possible outcomes

I assume that you are asking for the probability of getting a 1 or a 2 or a 3.

In this case, the number of desired outcomes is 3 (for values 1 or 2 or 3) and the number of total possible outcomes is 6.

So, the solution is 3/6 = 1/2.

Another approach, is to calculate the probability of each event and add them.
probability of getting 1 = probability of getting 2 = probability of getting 3 = 1/6.

1/6 + 1/6 + 1/6 = 3/6 = 1/2

Hope it helps!
=============
Go for 700+!!
https://mbajourney12.blocked/
Join the discussion

by crackgmat007 » Fri Oct 16, 2009 6:36 pm
uymba wrote:the general rule for probablity is:

number of desired outcomes
---------------------------------
number of total possible outcomes

I assume that you are asking for the probability of getting a 1 or a 2 or a 3.

In this case, the number of desired outcomes is 3 (for values 1 or 2 or 3) and the number of total possible outcomes is 6.

So, the solution is 3/6 = 1/2.

Another approach, is to calculate the probability of each event and add them.
probability of getting 1 = probability of getting 2 = probability of getting 3 = 1/6.

1/6 + 1/6 + 1/6 = 3/6 = 1/2

Hope it helps!
I guess question is 1, 2 AND 3. In this case, is 1/6 * 1/6 * 1/6 the answer?
Join the discussion

by sanjana » Mon Oct 19, 2009 9:48 pm
The Question doesnt say that u are replacing the number after picking,

So P(PICKING 1) = 1/6
P(picking 2) = 1/5
P(picking 3) = 1/4

P(picking 1,2 and 3) is 1/6*5*4 = 1/120

whats the OA?
Join the discussion